OpenStudy (anonymous):

What is the solution set of the following equation? 6x^2 = 13x + 5

OpenStudy (anonymous):

{-1/3, 5/2}? 2x + 5 = 0 2x = -5 x = -5/2 (-2.5) 3x - 1 = 0 3x = 1 x = 1/3

OpenStudy (anonymous):

oopsuies

OpenStudy (lgbasallote):

\[\implies 6x^2 -13x - 5 = 0\] \[\implies (3x+1)(2x-5)=0\]

OpenStudy (lgbasallote):

seems you made a mistake with signs

OpenStudy (anonymous):

yeah..sorry but seems like its right

OpenStudy (anonymous):

if i change the signs

OpenStudy (lgbasallote):

yep..change the signs and you're right already